The perimeter of a rectangle is 800 feet, and length of it is 50 feet less than half the width. What are it's dimensions?
Answer by checkley77(12844)

L=W/2-50
P=2L+2W
800=2(W/2-50)+2W
800=W-100+2W
800+100=3W
900=3W
900/3=W
3OO=W ANS. FOR THE WIDTH.
L=300/2-50
L=150-50
L=100 ANS. FOR THE LENGTH.
PROOF:
800=2*100+2*300
800=200+600
800=800